Show that the closed sphere with centre (1,3,5) and radius 8 in R³ is contained in the


open cube


P ={(x, y, z): | x -1| <10, | y - 3 | < 10, | z - 5 | <10}.


1
Expert's answer
2022-01-24T15:34:54-0500

The equation of the sphere with centre "(1, 3, 5)" and radius "8" in "\R^3" is



"(x-1)^2+(y-3)^2+(z-5)^2=8^2"

"(x-1)^2\geq0, x\in \R"

"(y-3)^2\geq0, y\in \R"

"(z-5)^2\geq0, z\in \R"

Then


"0\leq(x-1)^2\leq8^2""|x-1|\leq8""0\leq(y-3)^2\leq8^2""|y-3|\leq8""0\leq(z-5)^2\leq8^2""|z-5|\leq8"

Hence


"|x-1|<10, |y-3|<10, |z-5|<10, x,y,z\in \R"


This means that the closed sphere with centre "(1,3,5)" and radius "8" in "\R^3" is contained in the open cube

"P = \{(x, y, z): |x − 1| <10, |y − 3| <10, |z − 5| <10\}."